# QDsim: An user-friendly toolbox for simulating large-scale quantum dot device

### v1

@@ II.1 Derivation of the Constant-Capacitance Model Energy Equation
  • the dot-to-dot mutual capacitance matrix, in which each element $\widetilde{C^{DD}_{ij}}$ represents the capacitive coupling between dot $i$ and dot $j$, and the diagonal elements represent the dot’s self capacitance; and
@@
Assuming $\mathbf{Q}_{D}=e\mathbf{N}_{D}$, where $e$ is the elementary charge with sign ($+1$ for holes, $-1$ for electrons), we can rewrite the free energy function in terms of $\mathbf{N}_{D}$ (the charge configuration that we want to obtain) and $\mathbf{V}_{G}$ (the gate voltages that we want to tune). Omitting gate self-energies, we rewrite the free energy from Equation (7) as
$F\left(\mathbf{N}_{D};\mathbf{V}_{G}\right)=$ $\frac{1}{2}\left(e^{2}\mathbf{N}^{T}_{D}\mathbf{C}^{-1}_{DD}\mathbf{N}_{D}\right.$
$-2e\mathbf{N}^{T}_{D}\mathbf{C}^{-1}_{DD}\mathbf{C}_{DG}\mathbf{V}_{G}$
@@
For the remainder of the paper, we will adopt units such that $|e|=1$, and define the charging-energy matrix $\mathbf{E}_{C}=\mathbf{C}^{-1}_{DD}$. It is worth noting that we will need to take into account the sign of the charge, which will be negative for electrons, and positive for holes simulations.
@@ II.2 Ground States and Coulomb Polytopes in V-Space
Direct consequence of the inequality (11) is that the regions in $V_{G}$-space admitting a particular occupation $\mathbf{N}$ as a ground state form convex polytopes. We can also prove that two polytopes sharing an interior point $\mathbf{V}_{0}$ must coincide, implying that two states $\mathbf{N}_{1}$ and $\mathbf{N}_{2}$ are degenerate if and only if $\mathbf{N}_{1}-\mathbf{N}_{2}\in\mathrm{Null}(\mathbf{C}^{T}_{DG}\mathbf{E}_{C})$.
@@ III.1 The quantum dot device class: QDDevice
  • evaluate_dot_dot_mutual_capacitance_matrix: This method calculates the dot-to-dot mutual capacitance matrix of the QDDevice based on the dot locations. It employs a distance-based model to compute the capacitance interactions, taking into account the arrangement of quantum dots.
  • evaluate_dot_gate_mutual_capacitance_matrix: This method is used to compute the dot-to-gate mutual capacitance matrix based on the dot locations. It assumes that each gate corresponds to a dot in only controls that dot, making it suitable for certain device configurations.
@@ IV.1.4 Plotting and adding noise
The plot_charge_stability_diagrams method visualizes the simulation outcomes, plotting either the potential or current landscape, with an option to incorporate noise. To visualize the sensed potential, the plot_potential argument should be set to True. Conversely, setting plot_potential to False directs the method to plot the sensed current.
@@
Figure 3: Simulated charge stability diagrams for a double dot device. In Figure (a) the potential sensed is plotted without noise. In Figure (b) and (c) the gradient is evaluated, therefore plotting the current, with and without noise.
@@ IV.2.1 Simulation
Figure 5: Simulated charge stability diagrams for a 4x4 shared-control dot device. In Figure (a) the potential sensed is plotted without noise. In Figure (b) and (c) the gradient is evaluated, therefore plotting the current, with and without noise correspondingly.
@@ IV.3.1 Individual control: dot location specification
Figure 7: Simulated charge stability diagrams for a custom design with individual control. In Figure (a) the potential sensed is plotted without noise. In Figure (b) and (c) the gradient is evaluated, therefore plotting the current, respectively with and without noise.
@@ IV.3.2 Shared control: specifying the dot-to-gate mutual capacitance matrix
Figure 9: Simulated charge stability diagrams for a custom design with shared control. In Figure (a) the potential sensed is plotted without noise. In Figure (b) and (c) the gradient is evaluated, therefore plotting the current, respectively with and without noise.
@@ V Performance Evaluation and Constraints
Among comparable available tools, the QTT package [30] stands out, although it primarily addresses highly specialized experimental scenarios. QTT is geared more towards analysis and measurements within precise experimental frameworks, rendering it somewhat challenging for beginners and limiting its applicability to a narrow range of devices (e.g., double dots, in-line 6 dot array, triple dot, square dot) all under individual control schemes. This specificity in focus is the primary reason a detailed comparison with QTT has not been pursued, as the two packages cater to distinct needs and applications. Furthermore, very recently a new package named SimCATS [31] has been published. It aims to achieve a maximally realistic description of CSD especially with regards to sensing and noise implementations. While having a different focus, it caters to the need of realistic noise and could be a promising addition to our approximated solution. Furthermore, our work is complementary to a simultaneously published package called QDarts [32], which leverages a polytope-finding algorithm to efficiently simulate and locate charge transitions in the presence of tunnel couplings, non-constant charging energy and realistic noise. However due higher computational complexity, it focuses on smaller dot arrays (approximately 10 dots).
@@
When considering limitations, we identify two main categories: physical and functional constraints. On the physical side, it is crucial to acknowledge that the simulations are based on electrostatics and do not account for quantum mechanical effects. This limitation impacts the simulations’ realism, potentially affecting the precision in modeling actual devices. However, the package’s primary goal is not to perfectly replicate physical systems but rather to generate datasets for machine learning applications, where models can be initially trained on sufficiently accurate simulated data and later fine-tuned with real experimental data.
@@ VI Conclusion
The development of this package was driven by the ambition to offer the scientific community a highly accessible and user-friendly tool, specifically designed to streamline the generation of charge stability diagrams for extensive quantum dot arrays. With a focus on efficiency and speed, this package aims to significantly reduce the time and complexity traditionally associated with such simulations. By simplifying the process for both students and professionals, whether in theoretical or experimental domains, the package opens up new avenues for exploration and discovery in the field of quantum computing and nanotechnology. Furthermore, it lays the groundwork for the creation of comprehensive datasets, essential for the advancement of machine learning applications within this sphere.
@@ Acknowledgements
The authors acknowledge fruitful discussions with Francesco Borsoi, Brennan Undseth, Maia Rigot, and Menno Veldhorst.

### v2

# QDsim: A user-friendly toolbox for simulating large-scale quantum dot devices

@@ II.1 Derivation of the Constant-Capacitance Model Energy Equation
  • the dot-to-dot mutual capacitance matrix, in which each element $\widetilde{C^{DD}_{ij}}$ represents the capacitive coupling between dot $i$ and dot $j$, and the diagonal elements represent the dot’s self capacitance;
@@
Here $\mathbf{Q}_{D}=q\mathbf{N}_{D}$, where $q$ is the unit charge of the carriers. For electrons, $q=-e$ and for holes, $q=+e$, with $e$ being the absolute value of the elementary charge ($e=1.602\times 10^{-19}$ Coulombs). We can rewrite the free energy function in terms of $\mathbf{N}_{D}$ (the charge configuration that we want to obtain) and $\mathbf{V}_{G}$ (the gate voltages that we want to tune). Omitting gate self-energies, we rewrite the free energy from Equation (7) as
$F\left(\mathbf{N}_{D};\mathbf{V}_{G}\right)=$ $\frac{1}{2}\left(q^{2}\mathbf{N}^{T}_{D}\mathbf{C}^{-1}_{DD}\mathbf{N}_{D}\right.$
$-2q\mathbf{N}^{T}_{D}\mathbf{C}^{-1}_{DD}\mathbf{C}_{DG}\mathbf{V}_{G}$
@@
For the remainder of the paper, we will adopt units such that $|q|=e=1$, and define the charging-energy matrix $\mathbf{E}_{C}=\mathbf{C}^{-1}_{DD}$. It is worth noting that we will need to take into account the sign of the charge, which will be negative for electrons, and positive for holes simulations.
@@ II.2 Ground States and Coulomb Polytopes in V-Space
A direct consequence of the inequality (11) is that the regions in $V_{G}$-space admitting a particular occupation $\mathbf{N}$ as a ground state form convex polytopes. We can also prove that two polytopes sharing an interior point $\mathbf{V}_{0}$ must coincide, implying that two states $\mathbf{N}_{1}$ and $\mathbf{N}_{2}$ are degenerate if and only if $\mathbf{N}_{1}-\mathbf{N}_{2}\in\mathrm{Null}(\mathbf{C}^{T}_{DG}\mathbf{E}_{C})$.
@@ III.1 The quantum dot device class: QDDevice
  • evaluate_dot_dot_mutual_capacitance_matrix: This method calculates the dot-to-dot mutual capacitance matrix of the QDDevice based on the dot locations. It employs a distance-based model to compute the capacitance interactions, taking into account the arrangement of quantum dots. For each pair of dots, the mutual capacitance is determined based on their physical separation assuming electron-electron or hole-hole Coulomb interaction. The diagonal elements of the matrix represent the self-capacitance of each dot, which is initially set to a user-defined value (c0). Off-diagonal elements are calculated based on the inverse square of the distance between dots to approximate the capacitive coupling.
  • evaluate_dot_gate_mutual_capacitance_matrix: This method is used to compute the dot-to-gate mutual capacitance matrix based on the dot locations. It assumes that each gate corresponds to a dot and only controls that dot, making it suitable for certain device configurations. Similar to the dot-to-dot capacitance, for each dot-gate pair, the mutual capacitance is determined from the Coulomb interaction based on their physical separation (where the gate controlling the corresponding dot is considered as placed in the same coordinates of the dot it controls), with an additional factor to account for crosstalk effects between neighboring gates. Random noise can be added to simulate variations in gate capacitances and to introduce crosstalk. It is important to highlight that unlike [27] [28], our focus is not to have a precise description of the electrostatic environment. Our goal is to provide the user with a ready-to-use function that provides a good enough high-level description.
@@ IV.1.4 Plotting and adding noise
The plot_charge_stability_diagrams method visualizes the simulation outcomes, plotting either the potential or current landscape, with an option to incorporate noise. To visualize the sensed potential, the plot_potential argument should be set to True. Conversely, setting plot_potential to False directs the method to plot the sensed current. It must be noted that the current is not explicitly calculated based on physical tunnel couplings or temperature dependencies. Instead, we take the gradient of the sensed potential matrix to estimate the current. This approach is a simplified method to provide a qualitative visualization of how the sensed potential changes. This method does not account for tunnel couplings between dots, source and drain locations, or temperature effects.
@@
By incorporating noise into the simulated data, we aim to create more realistic training datasets for machine learning models. These models are often used for automating the tuning and control of quantum dot devices, and training them on noise-free data could lead to suboptimal performance in real experimental conditions. Furthermore, machine learning models trained on data with various types and levels of noise tend to be more robust and generalize better to unseen data [32].
@@
Figure 3: Simulated charge stability diagrams for a double dot device. (a) Potential sensed without noise. (b) Current calculated from the gradient of the potential matrix, without noise. (c) Current calculated from the gradient of the potential matrix, with noise.
@@ IV.2.1 Simulation
Figure 5: Simulated charge stability diagrams for a 4x4 shared-control dot device. (a) Potential sensed without noise. (b) Current calculated from the gradient of the potential matrix, without noise. (c) Current calculated from the gradient of the potential matrix, with noise.
@@ IV.3.1 Individual control: dot location specification
Figure 7: Simulated charge stability diagrams for a custom design with individual control. (a) Potential sensed without noise. (b) Current calculated from the gradient of the potential matrix, without noise. (c) Current calculated from the gradient of the potential matrix, with noise.
@@ IV.3.2 Shared control: specifying the dot-to-gate mutual capacitance matrix
Figure 9: Simulated charge stability diagrams for a custom design with shared control. (a) Potential sensed without noise. (b) Current calculated from the gradient of the potential matrix, without noise. (c) Current calculated from the gradient of the potential matrix, with noise.
IV.4 Running QDsim on a cluster
To enhance the utility of QDsim for generating datasets, we have implemented the QDParallelSimulator class. This class leverages the parallelization capabilities of the mpi4py package, allowing for efficient and scalable simulations. Users can define the devices and configurations they wish to simulate, as illustrated below:
Python Code
simsetups = [
    ’qd_device’: device,
    ’simulation_type’: ’Electrons’,
    ’configurations’: [
        # List of Tuples. Each Tuple
        represents the configuration
        for a single simulation of the
        charge stability diagram for
        the device.
        # Configurations are specified
        with a Tuple with the following
        elements:
        #   - sensor_location
        #   - scanning_gate_indexes
        #   - gates_voltages
        #   - fixed_voltage
        ((2, 1), (0, 1), None, 1.5),
        ((1, 2), (2, 3), (1.89, 1.56,
        None, None), None),
    ’n_points_per_axis’: 60,
    ’v_range_x’: [-5, 20],
    ’v_range_y’: [-5, 20],
    ’solver’: ’SCIP’,
    ’save_voltage_occupation_data’: True,
    ’save_sensing_data’: True,
    ’save_current_data’: True,
    ’save_charge_stability_diagram_plot’
        : True,
    ’charge_stability_diagram_plot_settings’
        ’cmapvalue’: ’RdPu’,
        ’gaussian_noise’: False,
        ’white_noise’: False,
        ’pink_noise’: False,
        ’plot_potential’: False,
        # Noise parameters
        ’gaussian_noise_params’: None,
        ’white_noise_params’: None,
        ’pink_noise_params’: None,
        ’plot_format_extension’: ’pdf’
After defining a quantum dot device, users can utilize a dictionary format to specify all the necessary attribute values for running the simulation. In the example provided, we simulate one device in two different configurations. The keyword ’configurations’ contains a list of two elements, each identifying a unique simulation configuration, including sensor locations, the gates to be scanned, and the plunger gate voltage values. The list simsetups defined above can contain any number of dictionaries, with each dictionary corresponding to a specific device being simulated. The code can thus be executed on any cluster, facilitating and accelerating the process of dataset creation.
@@ V Performance Evaluation and Constraints
Among comparable available tools, the QTT package [33] stands out, although it primarily addresses highly specialized experimental scenarios. QTT is geared more towards analysis and measurements within precise experimental frameworks, rendering it somewhat challenging for beginners and limiting its applicability to a narrow range of devices (e.g., double dots, in-line 6 dot array, triple dot, square dot) all under individual control schemes. This specificity in focus is the primary reason a detailed comparison with QTT has not been pursued, as the two packages cater to distinct needs and applications. Furthermore, very recently a new package named SimCATS [34] has been published. It aims to achieve a maximally realistic description of CSD especially with regards to sensing and noise implementations. While having a different focus, it caters to the need of realistic noise and could be a promising addition to our approximated solution. Additionally, another similar package, QArray [35], has been made available to the community. Like QDsim, it has a focus on simulation speed, while also including more effects such as latching and thermal broadening. It achieves much faster simulations results, in the order of milliseconds, for small-to-medium devices of up to 16 dots, but it falls short on simulating the large-scale devices with more than 50 dots, which has been the most successful goal of QDsim.
Furthermore, our work is complementary to a simultaneously published package called QDarts [36], which leverages a polytope-finding algorithm to efficiently simulate and locate charge transitions in the presence of tunnel couplings, non-constant charging energy and realistic noise. However due higher computational complexity, it focuses on smaller dot arrays (approximately 10 dots).
@@
Figure 11: Comparison of simulated results for three packages: (a) and (d) are from QDsim, (b) and (e) are from QArray, (c) and (f) are from QDarts. The column on the left, including figures (a), (b), and (c), shows the charge stability diagram of a double quantum dot device. Plot (c) is a close-up of an interdot transition. The column on the right, including figures (d), (e), and (f) shows the close up of the charge stability diagram of a quantum dot array of 6 dots. For the plots of QDsim and QArray, the same mutual capacitance matrices were used, while some random capacitance matrices where used in QDarts.
In addition, we compared QDsim with two other prominent simulation packages, QArray and QDarts. The comparative results are presented in Figure 11. This analysis evaluates the quality and features of our simulator relative to the other packages for two distinct device architectures: a double quantum dot (left column) and a six-dot array (right column).
The results demonstrate that all three simulators qualitatively capture the key high-level features of charge stability diagrams, including the interdot regions and clear charge transition lines. This consistency across the simulators is anticipated, as they are all based on the constant-interaction model.
However, QArray distinguishes itself by offering additional advanced features, such as the physical implementation of charge sensors, modeling of semi-classical phenomena like latching, thermal broadening, and various noise models. These enhancements contribute to a more realistic representation of the quantum dot behavior, as illustrated in Figure 11 (b) and (e). The inclusion of these features results in simulations that not only depict the fundamental charge configurations but also mimic the experimental conditions and phenomena, thereby providing a more comprehensive and realistic output.
When considering limitations, we identify two main categories: physical and functional constraints. On the physical side, it is crucial to acknowledge that the simulations are based on electrostatics and do not account for quantum mechanical effects. This limitation impacts the simulations’ realism, potentially affecting the precision in modeling actual devices. Beyond comparison to other effective models shown in Figure 11 there is a significant body of literature exploring full quantum simulation of quantum dot arrays based on Fermi-Hubbard model [37, 38, 39, 40, 41, 42]. Visual comparison of QDsim charge stability diagrams with these works confirm that constant interaction model is generally able to identify position and direction of charge transition without an ability to resolve more subtle features. In this context we want to emphasize that the primary goal of QDsim is not to perfectly replicate physical systems but rather to generate datasets for machine learning and other data driven applications, where models can be initially trained or test on sufficiently accurate simulated data and later fine-tuned with real experimental data. A recent example of this approach is illustrated in Ref. [43].
@@ VI Conclusion
The development of QDsim package was driven by the ambition to offer the scientific community a highly accessible and user-friendly tool, specifically designed to streamline the generation of charge stability diagrams for extensive quantum dot arrays. With a focus on efficiency and speed, this package aims to significantly reduce the time and complexity traditionally associated with such simulations. By simplifying the process for both students and professionals, whether in theoretical or experimental domains, the package opens up new avenues for exploration and discovery in the field of quantum computing and nanotechnology. Furthermore, it lays the groundwork for the creation of comprehensive datasets, essential for the advancement of machine learning applications within this sphere.
@@ Acknowledgements
The authors acknowledge fruitful discussions with Francesco Borsoi, Brennan Undseth, Maia Rigot, Sam Katiraee-Far and Menno Veldhorst and the help and collaboration of Barnaby van Straaten and Jan A. Krzywda in sharing their data with us. The project was supported by the Digital Competence Centre, Delft University of Technology.